A long jumper moving horizontally with a speed of 5.3 m/s has a kinetic energy of 235 J. Calculate the mass of the jumper.


1
Expert's answer
2021-12-09T09:12:06-0500

"KE=mv^2/2\to m=2KE/v^2=2\cdot235/5.3^2=16.7\ (kg)" . Answer